Prove (X - Y) nZ = (X n 2) - Y
Answer #1

ç - (x-2) - Y than from 0 (x-y) n2 (x-7) nz 4 ③ = (xnal - y Method we II know that x - у - хлу since (x-y) n z = (x nyo nz =